# Leveraging temporal features of the divergence measure to detect chaos in conservative systems

### v1

@@ 3.1. Time behavior of $\langle DIV\rangle$ in the original phase space
We computed the $DIV$ for a set of $200$ initial conditions distributed randomly into $0,2\pi]^{2}$ for parameters $K$ randomly chosen in the range $[0.6,4]$. By using the FLI computed at the final time $N$, each initial condition is assigned to a specific label (regular or chaotic). The Fig. 2 shows the time evolution of the divergences $DIVs$ for this set of initial conditions. Initial conditions that are FLI regular are color coded in blue, while FLI chaotic trajectories appear in red. We observe that the $DIVs$ tend to form two distinct clusters, indicating that the $DIV$ metric effectively captures the regular or chaotic nature of the orbit through its distinct time evolution. We note the spreading of the $DIVs$ values to be less pronounced in the case of regular orbits. The fits of the ensemble averages $\langle DIV\rangle$ of the regular and chaotic cluster lead to power laws (dashed black curves) with clearly distinct exponents. Regular orbits are characterised by
(8) $\langle DIV(N)\rangle\propto N^{\gamma},\,\gamma\sim-1,$
(numerically, we find $\gamma_{\textrm{reg}}=1.045$), whilst chaotic orbits experience a much slower decay rate (numerically, we estimate $\gamma_{\textrm{cht}}=-0.48$). It should be noted that the asymptotic decay of $\langle DIV\rangle$ in the regular case aligns with the asymptotic decay of the maximal Lyapunov characteristic exponent (Benettin et al.,, 1976; Contopoulos et al.,, 1978).
Figure 2. The measure $DIV$ as a function of the time series length, $N.$ The ensemble of 200 trajectories is separated according to the threshold values into regular (blue) and chaotic (red) parts, 152 and 48 members, respectively. The fit (dashed lines) to the ensemble averages (gray solids) shows different power-law exponents for the two kinds of motion. The green curve will be discussed thereafter.
@@
We further confirmed our observations by repeating the previous experiment for $4$ frozen values of $K$, namely $K\in\{0.6,1.1,2.6,4\}$. The chosen values are representative of phase spaces dominated by stability ($K=0.6$) or chaoticity ($K=4$). The intermediate value of $K=1.1$ reflects a mixed phase space regime, where chaotic and regular structures cohabit in a balanced proportion. Our previous observations fully extrapolate to these $4$ cases. In particular, regular trajectories have their $DIVs$ following closely the curve $N^{\gamma_{\textrm{reg}}}$, with $\gamma_{\textrm{reg}}\leq-1$, whilst chaotic trajectories have in average a slower decay rate, close to 1/2-1/2, as shown in Fig. 3. A finer analysis was performed as a function of the nonlinearity parameter $K$, although this required reducing the final time to $N=500$. We found $\gamma_{\textrm{cht}}\sim-1/2$ to be characteristic of the decay rate of $\langle DIV\rangle$ in the chaotic regime, see Fig. 4. (The power-laws with pronounced exponents seem to be characteristic of deterministic systems. For a stochastic ARMA process, the exponent is much smaller, reflecting the quasi absence of decay rate of $DIV$, see appendix A). In appendix B, we present numerical evidences that our results are robust to the parameters of the RPs, that is, when $RR$ varies in a realistic manner or when the norm to gauge the recurrence is changed. The appendix C extends our conclusions to two resonant archetypal $1$ degree-of-freedom non-autonomous Hamiltonian models where the divergence $DIV$, computed from the 2-dimensional discretisation of the dynamics via the stroboscopic map, also obeys the laws just described. Thus, there is a body of results that supports these numerical findings.
Figure 3. Same as in Fig. 2 for varying nonlinearity parameters $K$ ranging dynamical regimes dominated by stability ($K=0.6$) and chaoticity ($K=4$).
@@
Figure 4. Evolution of the power laws exponent $\gamma$ in regular or chaotic cases as a function of $K$ computed at $N=500$. The separation between regular and chaotic orbits is well reflected by the exponent. For regular components, the exponent is always smaller than $\gamma_{\textrm{reg.}}\leq 1$, whilst for chaotic ensembles, $\gamma_{\textrm{cht.}}$ is larger and tends to align to $\gamma_{\textrm{cht.}}=-1/2$ for very chaotic components.
A close inspection of specific curves in the set of chaotic trajectories of Fig. 2 reveals that some orbits have a final $DIV$ comparable to regular orbits. The green orbit of Fig. 2 is one of such orbit. We identified initial conditions and parameters of these orbits for a more detailed analysis and found that those orbits exhibit stickiness during their evolution. The sharp and successive drops in $DIV$ are correlated to time intervals during which the orbit is temporarily trapped in specific regions of the phase space. During this time, the decay of $DIV$ tends to align to the decay rate of regular trajectories, $\gamma_{\textrm{reg.}}=-1$. When the orbit leaves the sticky zone and experiences large excursion, the exponent is closer to $\gamma_{\textrm{cht.}}=-1/2$. Altogether, it produces the apparent staircase pattern. This dynamical behavior is also observed and confirmed through the FLI analysis, where the FLI behaves as plateau during during the sticky events (no hyperbolicity contributing to the growth of the norm of the tangent vector). The panel of Fig. 5 illustrates this phenomenology on a representative case. Thus, we find that the $DIV$ metric is sufficiently sensitive to detect the presence of sticky dynamics, which manifest as significant fluctuations in $DIV$ over specific time windows corresponding to temporary captures. (Note that this observation provides another RQA measure able to capture stickiness, besides the recurrence rate $RR$, see Palmero et al., (2022).)
Figure 5. (Left) Phase space of the chaotic orbit with initial condition marked in red extracted from Fig. 2. The orbit is chaotic with a low divergence value (green curve). Two portions of the orbit are colored (in blue and green) in accordance with the 22 most prominent sticky events encountered during its evolution. (Right, from top to bottom) Time evolution of the angle $\theta$, the FLI and the $DIV$. Sticky events (localised variations of $\theta$) correspond to plateau in the FLI and sharp decay in the DIVDIV measure, as visually materialised with the blue and green boxes serving as guides. The sharp decreases of DIVDIV during the sticky events explain the low final value of $DIV$, despite being chaotic.
The distinct power laws just revealed are based on the computation of the RP from the 2-dimensional trajectory. Adopting a nonlinear time series perspective, we now conduct similar numerical experiments starting from the knowledge of a specific univariate observable $z=g(p,\theta)$ over time, $\{z_{0},z_{1},\cdots,z_{N}\}$, $g$ being an observable function. The method of delay-coordinates often follows to reconstruct the phase-space. This is the approach taken subsection 3.3. For now, we compute the RP plot with no embedding, i.e.,without any phase space reconstruction. Given the low-dimensionality of the original system, according to Iwanski and Bradley, (1998), it is expected the RP plot to be almost independent from the reconstruction process. As the $DIV$ follows directly from RPs, we thus expect as byproduct the $DIV$ to be also almost independent from the latter. Under this setting, Eq. (1) becomes now
(9) $r_{i,j}=\Theta\big(|z_{i}-z_{j}|-\epsilon\big).$
Figure 6. Time evolution of $DIV$ as a function of time based on the two observables corresponding to $\theta$ (left) and momentum $p$ (right). No reconstruction of the phase space is performed.
Contrarily to the approach taken in subsection 3.2, starting from the univariate time series $\{z_{0},z_{1},\cdots,z_{N}\}$, we now reconstruct a phase space in $\mathbb{R}^{d}$, for some estimated $d$, via the delay-coordinates method with lag-time $\tau$ (Abarbanel et al.,, 1993). More specifically, let be $m=N-(d-1)\tau$, then the reconstructed trajectory reads $\{y_{i}\}_{i=0}^{m}$ with
(10) $y_{i}=(z_{i},z_{i+\tau},\cdots,z_{i+(d-1)\tau}).$
There are several methods to determine the embedding parameters $d$ and delay time $\tau$. Here, we fix the time delay to $\tau=1$ and determine $d$ using the false nearest neighbors method (Kantz and Schreiber,, 2003; Small,, 2005). Once the phase space is reconstructed, the divergence is computed from the RP matrix
(11) $r_{i,j}=\Theta\big(\left\lVert y_{i}-y_{j}\right\rVert-\epsilon\big),$
where $\left\lVert\bullet\right\rVert$ denotes the Euclidean norm in $\mathbb{R}^{d}$. Fig. 7 presents the analogue of Figs. 3 and 6 starting from the two observables $z(p,\theta)=\theta$ and $z(p,\theta)=p$ and extrapolates our previous observations. In appendix B, we present evidences that the results of the whole section do not chiefly depend on the choices made in computing the RPs: norm and threshold $\epsilon$.
Figure 7. Time evolution of $DIV$ as a function of time based on the two observables corresponding to $\theta$ (left) and momentum $p$ (right) in the reconstructed phase space.
[…]
To determine the threshold $\alpha$, we follow here standard approaches used for variational methods relying on the shape of the histogram of the metric computed for many initial conditions (Szezech Jr et al.,, 2005). Fig. 8 shows the phase space of the standard map for $K=1$ together with the landscapes of the FLI and $DIV$ for $250$ initial conditions spread along the dashed line joining the initial conditions $(\theta,p)=(\pi,0)$ and $(\pi,2\pi)$. For both indicators, we identify sharp increases, in 111-1 correspondence, when they cross transversely the thin chaotic layers. The bottom row of Fig. 8 shows the histograms of their final values for a resolved $400\times 400$ Cartesian mesh of initial conditions in $0,2\pi]^{2}$. The FLI distribution is rather right skewed, with the main peak located close to $\log(N)=\log(500)\sim 2.69$. This value characterizes the background of regular orbits. In this case, setting the threshold value $\alpha$ larger than this furnishes a reliable threshold to separate chaotic from regular trajectories, here set to $\alpha=4$. The distribution of the $DIVs$ is rather bimodal, where each mode reflects the typical values taken on regular and chaotic components. In this case, the threshold $\alpha$ is set as the minimal value between the two modes, leading here to $\alpha=2.15$. In the following, we systematically rely on histogram inspection to determine ad hoc $\alpha$ values (for regimes dominated by regularity, the histogram is skew right, for a regime dominated by chaoticity, it becomes skew left).
Fig. 9 portrays a series of dynamical maps corresponding to a scan of $0,2\pi]^{2}$ (top row) and a magnification of it (bottom row) materialised by the red square appearing near the $2:1$ periodic orbit. Each domain is meshed with a regular $400\times 400$ grid of initial conditions. Initial conditions leading to chaotic motions (either with the FLI or $DIV$) are marked in white, whilst regular orbits appear in black. Qualitatively, the dynamical maps of the first two columns of Fig. 9 are similar, meaning that the $DIV$ metric is indeed sound as chaos indicator. By plotting the initial conditions for which the labels are in disagreement, we are able to reveal the mismatch set (third column of Fig. 9). The two-scale analysis of the mismatch sets highlights that mismatch points are predominantly concentrated along the edges of the dynamical structures.
In order to compare the dynamical maps more quantitatively, we compute a simple metric with a straightforward interpretation: the probability of agreement $P_{A}$ between the labels produced (chaotic or regular) with either method over a specific domain $D$ (discretised as a regular Cartesian mesh of initial conditions.) The metric reads
(13) $P(D)=\frac{\#\textrm{Labels in agreement}}{\#\textrm{Initial conditions in $D$}}.$
Note that $P(D)$ concatenates the outputs of resolved dynamical maps. For $N=500$, we found the $DIV$ method to perform overall very well against the FLI, with agreement over $95\%$ over the range of nonlinearity parameters $[0.6,4]$, as illustrated in the left panel of Fig. 10. The range of values of $K\in[1,2.2]$ corresponds to a dip in $P(D)$. These dynamical regimes are characterized by regular and chaotic structures that cohabit in a rather balanced manner (mixed phase space regime). Given that mixed phase spaces contain many thin secondary structures, it is not surprising that the indicator is challenged. Nevertheless, the agreements are still excellent, above $95\%$. Increasing the time to $N=1,000$ doesn’t alter much this picture but improve very slightly the agreement. The right panel of Fig. 10 complements the former computation and shows, as a function of the nonlinearity parameter $K$, the volume of chaotic trajectories estimated either with the FLI or the $DIV$. The computation closely follows the one of Sander and Meiss, (2020). Throughout the range of $K$ values, we observe that the two curves closely follow one another, further demonstrating the effectiveness of the $DIV$ method in serving as a reliable chaos indicator.
Figure 10. (Left) Evolution of the probability of agreement $P_{A}$ (confer Eq. (13)) versus the nonlinearity parameter $K$. $P_{A}$ is computed over the domain $D=[0,2\pi]^{2}$. (Right) The size of the chaotic region in the phase space for different $K$ computed either with the FLI or the $DIV$. The two curves closely follow each other, strengthening further the use of $DIV$ as chaos indicator. Both panels are based on a mesh of $400\times 400$ initial conditions computed at $N=500.$
5. Conclusions
Although RPs and RQAs have been predominantly applied to dissipative systems, our study demonstrates the effectiveness and potential of the $DIV$ metric to serve as a simple diagnostic to detect chaos in conservative systems. We have compared the performance of this indicator against the fast Lyapunov indicator, a well established variational chaos detection method. We focused on the standard map, a paradigmatic discrete model of Hamiltonian chaos. From our extensive numerical simulations, we have quantitatively assessed the overall good agreement between the methods, further strengthening, but this time in the conservative regime, the strong correlation between the divergence measure and the presence of chaos in the dynamics. Our comparison has focused on moderately long orbits with $500$ data points, stepping towards real-world applicability of the method. The distinct power-laws of $DIV$ we have revealed in average depending on the regularity or chaoticity of the trajectories, on a much longer timescale, valid in the original and reconstructed phase space, shed also more light on asymptotic properties of the $DIV$ metric. In particular, we have observed a decay of $DIV$ as $1/N$ for regular orbits, which is interestingly the same rate as the maximal Lyapunov characteristic exponent. These distinct power laws could also be particularly valuable when dealing with a limited number of orbits, a context in which setting the threshold $\alpha$ from the histogram becomes challenging. Those properties suggest a possible new approach for chaos detection in time series, based on analyzing the slope of the decay rate of $DIV$ computed over windowed segments of the original series. Our current efforts are focused on leveraging this property, besides the noise free case, alongside providing analytical insights to support our findings.
Acknowledgments
This work was supported by the Hungarian National Research, Development and Innovation Office, under Grant No. K-153324 and TKP2021-NKTA-64 (T.K.), financed by the Ministry of Culture and Innovation of Hungary.
Data Availability Statement
The data presented in this study can be reproduced based on the models and list of parameters described in the text.
Appendix A $DIV$ versus $N$ for ARMA process
Appendix B Robustness with respect to choices
The power-laws and probability of agreements presented in the manuscript depend, a priori, on the choice of the norm and the neighboring condition defined by $\epsilon$ (fixed through a fixed $RR$.) Tables 1 and 2 show that the exponents we found are unaffected when the parameters vary. The boldfaced fonts repeat the exponents found in the manuscript’s corpus, presented here again for the sake of completeness. Only the case of the initial 22D trajectory has been explored.

### v2

@@ 3.1. Time behavior of $\langle DIV\rangle$ in the original phase space
We computed the $DIV$ for a set of $200$ initial conditions distributed randomly into the domain $0,2\pi]^{2}$ for parameters $K$ also randomly chosen in the range $[0.6,4]$. By using the FLI computed at the final time $N$, each initial condition is assigned to a specific label (regular or chaotic). The Fig. 3 shows the time evolution of the divergences $DIVs$ for this set of initial conditions. Initial conditions that are FLI regular are color coded in blue, while FLI chaotic trajectories appear in red. We observe that the $DIVs$ tend to form two distinct clusters, indicating that the $DIV$ quantifier effectively captures the regular or chaotic nature of the orbit through its distinct time evolution. We note the spreading of the $DIVs$ values to be less pronounced in the case of regular orbits.
The spatial averages $\langle DIV\rangle$ of the divergences on the regular and chaotic components are the central objects of the remainder of this section. They appear a solid bold lines in Fig. 3. Interestingly enough they both posses clear linear trends. Their best fits, determined with a Markov Chain Monte Carlo framework, lead indeed to power laws with well separated exponents (the posterior medians). Regular orbits are characterised by
(9) $\langle DIV(N)\rangle_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textrm{reg.}}}\propto N^{\gamma_{\textrm{reg.}}},\,\gamma_{\textrm{reg.}}\sim-1,$
numerically, we find $\gamma_{\textrm{reg}}=-1.044$. Chaotic orbits experience a much slower decay rate,
(10) $\langle DIV(N)\rangle_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textrm{cht.}}}\propto N^{\gamma_{\textrm{cht.}}},\,\gamma_{\textrm{cht.}}\sim-1/2,$
numerically, we find $\gamma_{\textrm{cht}}=-0.44$. The 16th and 84th percentiles of the samples all show a narrow gap. In both cases, the $R^{2}$ coefficients between the space averages and the power laws are $R^{2}=0.99$. We underline that the asymptotic decay of $\langle DIV\rangle_{\textrm{reg.}}$ echoes the asymptotic decay of the maximal Lyapunov characteristic exponent of regular trajectories (gBe76; gCo78).
Figure 3. Time evolution of the divergence DIVDIV as a function of the time-series length NN for an ensemble of $200$ trajectories of the standard map with varying $K\in[0.6,4]$. Regular orbits are shown in blue and chaotic trajectories in red. The spatial averages $\langle DIV\rangle$ over the regular and chaotic components (bold blue and red lines, respectively) exhibits clear power law behaviors, whose exponents are estimated numerically. The green curve corresponds to the time evolution of the divergence of a sticky orbit and will be discussed later.
@@
We further confirmed our observations by repeating the previous experiment for $4$ frozen values of $K$, namely $K\in\{0.6,1.1,2.6,4\}$. The chosen values are representative of phase spaces dominated by stability ($K=0.6$) or chaoticity ($K=4$). The intermediate value of $K=1.1$ reflects a mixed phase space regime, where chaotic and regular structures cohabit in a balanced proportion. Our previous observations fully extrapolate to these $4$ cases. In particular, regular trajectories have their $DIVs$ following closely the curve $N^{\gamma_{\textrm{reg}}}$, with $\gamma_{\textrm{reg}}\leq-1$, whilst chaotic trajectories have in average a slower decay rate, close to 1/2-1/2, as shown in Fig. 4 and Fig. 5.
Figure 4. Same as in Fig. 3 for frozen parameters $K$.
Figure 5. Same as in Fig. 3 for frozen parameters $K$.
A finer analysis was performed as a function of the nonlinearity parameter $K$, although this required, for computational convenience, reducing the final time to $N=500$. We found $\gamma_{\textrm{cht}}\sim-1/2$ to be characteristic of the decay rate of $\langle DIV\rangle$ in the chaotic regime, see Fig. 6. (The power laws with pronounced exponents seem to be characteristic of deterministic systems. For a stochastic ARMA process, the exponent is much smaller, reflecting the quasi absence of decay rate of $DIV$, see appendix A). We also underline that the 16th and 84th percentiles, on the chaotic component, tend to be smaller the more chaotic is the system (the larger the $K$). On the regular component, the percentiles are rather stable across the whole range of perturbing values.
@@
Figure 6. Evolution of the power laws exponent $\gamma$ on the regular and chaotic components as a function of $K$ computed at $N=500$. The separation between regular and chaotic orbits is well reflected by the different exponents, here shown with their 16th and 84th percentiles (vertical bars). For regular components, the exponent is most of the time smaller than $\gamma_{\textrm{reg.}}\leq 1$, whilst for chaotic components, $\gamma_{\textrm{cht.}}$ is much larger and tends to align to $\gamma_{\textrm{cht.}}=-1/2$. The two dashed lines, $\gamma=-1$ and $\gamma=-1/2$ are plotted as visual guides.
One may legitimately question the sensitivity of the obtained exponents with respect to the choice of parameters of the RPs. The tables 1 and 2 present the exponents we found by changing the norms (from $L_{2}$ to $L_{1}$ and $L_{\infty}$ norms) and the recurrence rates (decreasing the nominal value $RR=0.05$ to $RR=0.025$ and increasing it to $RR=0.1$). The close similarity of the exponents derived strengthens confidence in the robustness of the power laws with respect to the choice of RPs parameters.
@@
A close inspection of specific curves in the set of chaotic trajectories of Fig. 3 reveals that some orbits have a final $DIV$ comparable to regular orbits. The green orbit of Fig. 3 is one of such orbit. We identified initial conditions and parameters of these orbits for a more detailed analysis and found that those orbits exhibit stickiness during their evolution. The sharp and successive drops in $DIV$ are correlated to time intervals during which the orbit is temporarily trapped in specific regions of the phase space. During this time, the decay of $DIV$ tends to align to the decay rate of regular trajectories, $\gamma_{\textrm{reg.}}=-1$. When the orbit leaves the sticky zone and experiences large excursion, the exponent is closer to $\gamma_{\textrm{cht.}}=-1/2$. Altogether, it produces the apparent staircase pattern. This dynamical behavior is also observed and confirmed through the FLI analysis, where the FLI behaves as plateau during during the sticky events (no hyperbolicity contributing to the growth of the norm of the tangent vector). The panel of Fig. 7 illustrates this phenomenology on a representative case. Thus, we find that the $DIV$ quantifier is sufficiently sensitive to detect the presence of sticky dynamics, which manifest as significant fluctuations in $DIV$ over specific time windows corresponding to temporary captures. (Note that this observation provides another RQA measure able to capture stickiness, besides the recurrence rate $RR$, see mPa22.)
Figure 7. (Left) Phase space of the green trajectory of the sticky orbit of Fig. 3. Two portions of the orbit are colored (in blue and green) in accordance with the 22 most prominent sticky events encountered during its evolution. (Right, from top to bottom) Time evolution of the angle $\theta$, the FLI and the $DIV$. Sticky events (localised variations of $\theta$) correspond to plateau in the FLI and sharp decay in the DIVDIV measure, as visually materialised with the blue and green boxes serving as visual guides. The sharp decreases of DIVDIV during the sticky events explain the low final value of $DIV$, despite being chaotic, and the staircase pattern.
The distinct power laws just revealed are based on the computation of the RP from the two-dimensional trajectory. Adopting a nonlinear time series perspective, we now conduct similar numerical experiments starting from the knowledge of a specific univariate observable $z=g(p,\theta)$ over time, $\{z_{0},z_{1},\cdots,z_{N}\}$, $g$ being an observable function. The method of delay-coordinates often follows to reconstruct the phase-space. This is the approach taken subsection 3.3. For now, we compute the RP plot with no embedding, i.e.,without any phase space reconstruction. Given the low-dimensionality of the original system, according to jIw98, it is expected the RP plot to be almost independent from the reconstruction process. As the $DIV$ follows directly from RPs, we thus expect as byproduct the $DIV$ to be also almost independent from the latter. Under this setting, Eq. (1) becomes now
(11) $r_{i,j}=\Theta\big({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon_{i}}-|z_{i}-z_{j}|\big).$
Figure 8. Same as in Fig. 3 based on the two observables corresponding to $\theta$ (top) and momentum $p$ (bottom). No reconstruction of the phase space is performed.
Contrarily to the approach taken in subsection 3.2, starting from the univariate time series $\{z_{0},z_{1},\cdots,z_{N}\}$, we now reconstruct a phase space in $\mathbb{R}^{d}$, for some estimated $d$, via the delay-coordinates method with lag-time $\tau$ (hAb93). More specifically, let be $m=N-(d-1)\tau$, then the reconstructed trajectory reads $\{y_{i}\}_{i=0}^{m}$ with
(12) $y_{i}=(z_{i},z_{i+\tau},\cdots,z_{i+(d-1)\tau}).$
There are several methods to determine the embedding parameters $d$ and delay time $\tau$. Here, we fix the time delay to $\tau=1$ and determine $d$ using the false nearest neighbors method (hKa03; mSm05). Once the phase space is reconstructed, the divergence is computed from the RP matrix
(13) $r_{i,j}=\Theta\big({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\epsilon_{i}}-\left\lVert y_{i}-y_{j}\right\rVert\big),$
where $\left\lVert\bullet\right\rVert$ denotes the Euclidean norm in $\mathbb{R}^{d}$. Fig. 9 presents the analogue of Figs. 3, 4, 5 and 8 starting from the two observables $z(p,\theta)=\theta$ and $z(p,\theta)=p$ and extrapolates our previous observations.
3.4. Comments on the generality of the results
In subsection 3.1, we have revealed power laws with distinct exponents for the divergence quantifier on regular and chaotic components. We presented numerical evidences that our results do not chiefly depend on the parameters of the RPs (norms and recurrence rates). In subsections 3.2 and 3.3, we extended our conclusions starting from observables, and found similar results in reconstructed phase spaces. All these results are based on the standard map, a mapping derived from a Poincaré map of a continuous time dependent system. In Appendix B, we have considered Poincaré sections of two archetypal non-integrable resonant Hamiltonian models where resonances overlap, with one degree-of-freedom and non-autonomous. Our results and observations extrapolate to those models too, bringing further confidence to the generality of the observations made on the standard map model.
Figure 9. Same as in Fig. 3 based on the two observables corresponding to $\theta$ (top) and momentum $p$ (bottom) in the reconstructed phase space.
[…]
To determine the threshold $\alpha$, we follow here standard approaches used for variational methods relying on the shape of the histogram of the quantifier computed for many initial conditions. (This strategy is inspired by jSz05 in the context of finite-time Lyapunov exponent, and also used more recently in mrS22; mHi22; aCa25 with the Birkhoff average method or Lagrangian descriptors.) This is also the methodology we follow to binarise the FLI outputs.
Fig. 10 shows the phase space of the standard map for $K=1$ together with the landscapes of the FLI and $DIV$ for $250$ initial conditions spread along the dashed line joining the initial conditions $(\theta,p)=(\pi,0)$ and $(\pi,2\pi)$. For both indicators, we identify sharp increases, in 111-1 correspondence, when they cross transversely the thin chaotic layers. The bottom row of Fig. 10 shows the histograms of their final values for a resolved $400\times 400$ Cartesian mesh of initial conditions in $0,2\pi]^{2}$. The FLI distribution is rather right skewed, with the main peak located close to $\log(N)=\log(500)\sim 2.69$. This value characterizes the background of regular orbits. In this case, setting the threshold value $\alpha$ larger than this furnishes a reliable threshold to separate chaotic from regular trajectories, here set to $\alpha=4$. The distribution of the $DIVs$ is rather bimodal, where each mode reflects the typical values taken on regular and chaotic components. In this case, the threshold $\alpha$ is set as the minimal value between the two modes, leading here to $\alpha=2.15$. In the following, we systematically rely on histogram inspection to determine ad hoc $\alpha$ values (for regimes dominated by regularity, the histogram is skew right, for a regime dominated by chaoticity, it becomes skew left).
Fig. 11 portrays a series of dynamical maps corresponding to a scan of $0,2\pi]^{2}$ (top row) and a magnification of it (bottom row) materialised by the red square appearing near the $2:1$ periodic orbit. Each domain is meshed with a regular $400\times 400$ grid of initial conditions. Initial conditions leading to chaotic motions (either with the FLI or $DIV$) are marked in white, whilst regular orbits appear in black. Qualitatively, we found the dynamical maps of the first two columns of Fig. 11 to be very similar. These results support the validity of the $DIV$ quantifier as an indicator of chaos. By plotting the initial conditions for which the labels are in disagreement, we are able to reveal the mismatch set (third column of Fig. 11). The two-scale analysis of the mismatch sets highlights that mismatch points are predominantly concentrated along the edges of the dynamical structures.
In order to compare the dynamical maps more quantitatively, we compute a simple metric with a straightforward interpretation: the probability of agreement $P_{A}$ between the labels produced (chaotic or regular) with either method over a specific domain $A$ (discretised as a regular Cartesian mesh of initial conditions.) The metric reads
(15) $P_{A}=\frac{\#\textrm{Labels in agreement}}{\#\textrm{Initial conditions in $A$}}.$
Note that the scalar $P_{A}$ concatenates the outputs of resolved dynamical maps computed over the domain $A$. For $N=500$, we found the $DIV$ quantifier to perform overall very well against the FLI, with a probability of agreement over $95\%$ over the range of nonlinearity parameters $[0.6,4]$, as illustrated in the left panel of Fig. 12. For $K<1$, the probability of agreement is over $98\%$, just as for the values of $K>2.5$.
Figure 12. (Left) Evolution of the probability of agreement $P_{A}$ (confer Eq. (15)) versus the nonlinearity parameter $K$. $P_{A}$ is computed over the domain $A=[0,2\pi]^{2}$. (Right) The size of the chaotic region in the phase space for different $K$ computed either with the FLI or the DIVDIV on the same domain $A$. The two curves closely follow each other, strengthening further the use of $DIV$ as chaos indicator. Both panels are based on a mesh of $400\times 400$ initial conditions computed at $N=500.$
For the values of $K\in[1,2.2]$, a slight decrease in the agreement probabilities between the indicators is observed. Nevertheless, the agreements are still above $95\%$. These dynamical regimes are characterized by regular and chaotic structures that cohabit in a rather balanced manner (mixed phase space regime). The presence of numerous thin secondary structures naturally challenges the performance of the indicator, and possibly explain the dip. Extending the final time to $N=1,000$ does not markedly affect this picture, though it yields a slight improvement in the agreement. To further complement the results, the right panel of Fig. 10 shows, as a function of the nonlinearity parameter $K$, the volume of chaotic trajectories estimated either with the FLI or the $DIV$. The computation closely follows the steps of eSa20 with the weighted Birkhoff average. Throughout the range of nonlinear parameters, we observe that the two curves closely follow one another, further demonstrating the effectiveness of the $DIV$ method in serving as a reliable chaos indicator.
5. Conclusions
Although RPs and RQAs have been predominantly applied to dissipative systems, our study demonstrates the effectiveness and potential of the $DIV$ quantifier to serve as a simple diagnostic to detect chaos in conservative systems. We have compared the performance of this indicator against the fast Lyapunov indicator, a well established variational chaos detection method. We focused on the standard map, a paradigmatic discrete model of Hamiltonian chaos. From our extensive numerical simulations, we have quantitatively assessed the overall good agreement between the methods, further strengthening, but this time in the conservative regime, the strong correlation between the divergence measure and the presence of chaos in the dynamics. Our comparison has focused on moderately long orbits with $500$ data points, stepping towards real-world applicability of the method. The distinct power laws of $DIV$ we have revealed in average depending on the regularity or chaoticity of the trajectories, on a much longer timescale, valid in the original and reconstructed phase space, shed also more light on asymptotic properties of the $DIV$ quantifier. In particular, we have observed a decay of $DIV$ as $1/N$ for regular orbits, which is interestingly the same rate as the maximal Lyapunov characteristic exponent. On chaotic components, a decay rate closer to $1/\sqrt{N}$ is usually observed. These distinct power laws could also be particularly valuable when dealing with a limited number of orbits, a context in which setting the threshold $\alpha$ from the histogram becomes challenging. Those properties suggest a possible new approach for chaos detection in time series, based on analyzing the slope of the decay rate of $DIV$ computed over windowed segments of the original series. Our current efforts are focused on leveraging this property, besides the noise free case, alongside providing analytical insights to support our findings.
Acknowledgments
This work was supported by the Hungarian National Research, Development and Innovation Office, under Grant No. K-153324, K-152888, and TKP2021-NKTA-64 (T.K.), financed by the Ministry of Culture and Innovation of Hungary.
Data Availability Statement
The data presented in this study can be reproduced based on the computational notebooks availabe on Zenodo: https://doi.org/10.5281/zenodo.18356624 (tKo26).
Appendix A $DIV$ versus $N$ for ARMA process